OpenStudy (james1769):

Solve x2 + 8x − 3 = 0 using the completing-the-square method. (2 points) x = four plus or minus the square root of three x = negative four plus or minus the square root of three x = four plus or minus the square root of nineteen x = negative four plus or minus the square root of nineteen

OpenStudy (freckles):

do you know what value of k we can choose so the following is true: \[x^2+8x+?^2=(x+?)^2\]

OpenStudy (freckles):

value of ?*

OpenStudy (freckles):

hint: recall the following: \[x^2+bx+(\frac{b}{2})^2=(x+\frac{b}{2})^2\]

OpenStudy (james1769):

ok

OpenStudy (james1769):

im still confused

OpenStudy (freckles):

examples: \[x^2+2x+1=(x+1)^2 \\ x^2+4x+4=(x+2)^2 \\ x^2+6x+9=(x+3)^2 \\ x^2+8x+?^2=(x+?)^2\]

OpenStudy (freckles):

notice the 3rd term on the left hand side is just the number in front of x divided by two then squared this is why we have: \[x^2+bx+(\frac{b}{2})^2=(x+\frac{b}{2})^2\]

OpenStudy (james1769):

but which one is it then??

OpenStudy (james1769):

@imqwerty can u help me?

OpenStudy (freckles):

8 divided by 2 is 4...

OpenStudy (freckles):

\[x^2+bx+(\frac{b}{2})^2=(x+\frac{b}{2})^2 \\ x^2+8x+(\frac{8}{2})^2=(x+\frac{8}{2})^2 \\ \] so you have after simplifying: \[x^2+8x+4^2=(x+4)^2 \\ \text{ or } \\ x^2+8x+16=(x+4)^2 \]

OpenStudy (freckles):

\[x^2+8x-3=0 \\ x^2+8x=3 \\ x^2+8x+\text{ ___ }=3+ \text{___}\]

OpenStudy (freckles):

so what do you need to add on both sides to complete the square on the left hand side

OpenStudy (james1769):

sqr root

OpenStudy (freckles):

\[x^2+8x+16 \text{ is } (x+4)^2 \\ x^2+8x+ \text{___}=3+\text{___}\] so you need to add 16 on both sides in order to complete the square on the left hand side

OpenStudy (freckles):

did you want to tell me how far you gotten?

OpenStudy (freckles):

I could check your work so far.

OpenStudy (freckles):

well if get back my notifications aren't working so try mentioning me